OpenStudy (anonymous):

Solve Shane invested $3,500 in a savings account that pays 4% interest semiannually (every 6 months). How much money will he have in his account in 5 years

OpenStudy (anonymous):

you would use y=a(1+R)^x a =total amount of money r= rate ( to find r, move the decimal place over two places or divide by 100 x= time so fill it in a= 3,500 r=0.04 x= 5 years so y= 3,500(1+0.04)^5 you do the math, & you'll get your answer (: